OpenStudy (buckybarnes):

Check my answer. ∠1 and ∠2 are supplementary angles. M∠1is 4y+7 and m∠2 is 9y+4. Find m∠2

OpenStudy (buckybarnes):

I think m ∠2 is 121

OpenStudy (buckybarnes):

@3mar

OpenStudy (3mar):

We know that \[ang 1+ang 2=180\] then \[(4y+7)+(9y+4)=180\] simplify \[13y+11=180\] \[13y=169\] \[y=\frac{ 169 }{ 13 }=13\] substitute in ang 2: \[ang 2=(9y+4)=9(13)+4=121^o\] Hope that helps

OpenStudy (3mar):

I am going to pray now. Any thing or comment, just leave it and I reply when I come back. Salam.

OpenStudy (buckybarnes):

So it is 121?

OpenStudy (3mar):

Yes, it is 121. Good.

OpenStudy (3mar):

Thanks for the medal!